# A Tri-Modal Dataset and a Baseline System for Tracking Unmanned Aerial Vehicles

### v1

@@ II-B UAV-Targeted Tracking Benchmarks
For single-object tracking, Chen et al. [USC-Drone] introduced USC Drone, a dataset with 60 sequences and 20K frames, accompanied by a Fast-RCNN + MDNet system where the residual image input improves adaptation to a fast UAV motion. Huang et al. [Anti-UAV410] constructed Anti-UAV410, a large-scale thermal infrared benchmark with 410 sequences and 438K frames, and proposed SiamDT with dual-semantic features for tiny targets. Zhu et al. [Anti-UAV600] released Anti-UAV600, comprising 600 thermal sequences and 723K frames without prior bounding-box initialisation, and designed a “global detection + local tracking” framework that avoids reliance on the initial frame information through evidential collaboration. Despite the progress, these benchmarks remain limited to SOT and cannot depict scenarios involving multiple UAVs.
For multi-UAV tracking, UAVSwarm [UAVSwarm] builds a visible-light dataset with 72 sequences (12K frames) and validates a Faster-RCNN/YOLOX + ByteTrack baseline. MOT-FLY [MOT-FLY] presents an air-to-air high-resolution dataset (1920×1080) with 16 sequences (11K frames) and benchmarked advanced MOT frameworks. Recently, the 4th Anti-UAV Challenge [4th-Anti-UAV-workshop] released the first large-scale infrared multi-UAV benchmark with 300 sequences and 225K frames.
@@ II-C Multi-Modal Tracking Potential
The performance bottleneck exhibited by a single visual modality in complex scenarios have driven the development of multi-modal fusion technologies, such as RGB-T (RGB + thermal infrared), RGB-D (RGB + depth), and RGB-E (RGB + event). Integrating the complementary information conveyed by different modalities significantly enhances the tracking robustness.
@@
However, multi-modal research in anti-UAV tasks remains scarce. The exceptions include the Anti-UAV dataset [Anti-UAV], which was the first to introduce an RGB-T dataset, with 348 sequences and 297K frames. As a baseline, [Anti-UAV] proposed a dual-stream feature semantic consistency training strategy. Yet, due to the modalities being unaligned, it cannot exploit the real potential of the cross-modal complementarity, instead using them independently. Similarly, the Halmstad Drone dataset [Halmstad-Drone] includes 650 videos (365 for IR and 285 for RGB) along with 90 audio clips, but focuses on building automated multi-sensor systems rather than investigating deep learning-based fusion. Its incomplete correspondence between IR and visible-light data also limits the utility of multi-modal anti-UAV research. Moreover, these datasets primarily target detection or SOT, leaving a void in multi-modal multi-UAV tracking benchmarks and methods.
@@
A fundamental challenge in multi-modal tracking lies in modality misalignment. While most studies assume the data being spatio-temporally aligned, real-world data often suffers from misalignment due to differences in imaging principles, field-of-view, resolution, and transmission delays. This problem is especially critical for UAV tracking, where the targets are small and fast-moving, and even slight offsets undermine effective fusion. Traditional pixel-level solutions rely on feature-point matching and affine/perspective transformations to achieve alignment, but these methods incur considerable preprocessing latency and are unsuitable for real-time multi-UAV tracking. Recent studies, therefore, emphasise learnable feature alignment. Zhang et al.[ARCNN] proposed an aligned region CNN to align RGB and thermal regions using a Region Feature Alignment Module with RoI jitter. Chen et al.[CSOM] proposed cross-modality spatial offset modelling, together with an offset-guided deformable alignment and fusion, achieving robust UAV RGB-IR detection without strict pre-alignment. Inspired by these advances, we adopt two feature-level alignment strategies for UAV tracking: (1) learning sampling offsets via deformable convolution [def-conv]; and (2) predicting affine transformation matrices via a spatial transformer network [STN]. Both strategies are end-to-end trainable without additional losses, enabling efficient mitigation of misalignment in multi-modal fusion.
@@
Additionally, given the small size and irregular motion of UAVs, as well as the inability of traditional RGB or IR modalities to distinguish targets from background under similar appearance or temperature, this study further introduces the Event modality. As a cutting-edge visual sensor, event cameras capture asynchronous event streams by recording changes in light intensity. They offer unique advantages such as high dynamic range and microsecond-level temporal resolution, making them particularly suited for extreme scenarios. Existing event-based tracking research has produced datasets such as VisEvent [visevent], FE108 [fe108], EventVOT [eventvot], and COESOT [ceutrack], along with methods like CEUTrack [ceutrack], TENet [TENet], ViPT [vipt], and AFNet [afnet], but are restricted to the single-object tracking task, leaving a critical gap in event-based multi-UAV tracking.
@@ III MM-UAV BENCHMARK
The absence of a dedicated multi-modal multi-UAV tracking dataset to support anti-UAV research has significantly impeded the study of multi-modal fusion technologies. To address this limitation, this paper presents the first large-scale tri-modal multi-UAV tracking dataset, with each modality containing 1321 sequences (2.8 million frames), aiming to fill this critical research void. This section gives a comprehensive overview of the dataset, covering core aspects such as data acquisition, annotation protocols, statistical distributions, and challenging attributes.
@@ IV-D MMA-SORT
Existing multi-object tracking models (such as SORT [sort] and its variants [DeepSORT, deep-ocsort, ocsort, bytetrack, botsort, boostTrack, boostTrack++]) exhibit notable limitations in UAV-target tracking tasks. First, the small size and complex motion patterns of UAV targets pose a significant challenge to motion modelling. The linear motion assumption inherent to the Kalman filter often leads to substantial discrepancies between the predicted bounding boxes (based on previous frames) and the actual detections in the current frame. This renders trackers centred on IOU metrics ineffective in achieving cross-frame identity association, thereby resulting in a large number of ID switches (IDSW). To address this issue, we downplay the role of IOU metrics in the process of target identity association. Unlike traditional trackers that treat IOU matching as an essential prerequisite for identity association, we argue that Kalman filters are less reliable in modelling the motion of small, irregularly moving UAV targets, compared to targets like pedestrians and vehicles. Thus, rigidly adhering to IOU matching as a prerequisite in multi-UAV tracking tasks is overly restrictive.
@@
where $\mathbf{d}$ represents the final distance, $d_{\text{iou}}$ is the IOU distance between the detection box and the trajectory’s predicted box, and $d_{\text{app}}$ is the appearance embedding distance between the detection and the trajectory. The calculation for IOU distance and appearance distance follows those in BoT-SORT [botsort].
@@ V-A Implementation Details
In the proposed framework, the detector’s backbone, the FPN module, and the detection head follow the original YOLOX [YOLOX]. For RGB images (640×360) and IR images (640×512) from the two modalities, we first resize them to 640×640 using the standard YOLOX preprocessing pipeline before feeding them into the backbone for feature extraction. The proposed OGAA and ADFM modules are inserted between the backbone and FPN. They are responsible for cross-modal alignment and fusion of the three-scale feature maps extracted independently by the two modality-specific backbones. The fused multi-scale features are then forwarded to the respective FPN modules for subsequent detection, and the final detection results are passed to the MMA-SORT tracker for multi-object tracking.
@@
For the appearance-based ReID module, given the limited visual information of UAV targets, we directly adopt the ReID model and training strategy of DeepSORT [DeepSORT], using a ResNet backbone to extract discriminative appearance features.
During inference, the detection confidence threshold is set to 0.3. For the tracker, the IoU threshold $\theta_{\text{iou}}$ is set to 0.9, the appearance threshold to 0.2, the motion threshold $\theta_{\text{motion}}$ to 0.2, and the confidence threshold for initialising new tracks to 0.7.
@@ V-B Quantitative Comparison
For comparative evaluation, two major categories of state-of-the-art multi-object tracking methods were systematically evaluated, including Tracking by Detection approaches (DeepSORT [DeepSORT], BoT-SORT [botsort], OC-SORT [ocsort], Deep OC-SORT [deep-ocsort], BoostTrack [boostTrack], BoostTrack++ [boostTrack++]) and Tracking by Query approaches (TransTrack [transtrack], MOTR [motr], MOTRv2 [motrv2], MeMOTR [memotr], MOTIP [motip]). All trackers are retrained and tested on our MM-UAV to enable a fair comparison. Since existing methods are designed for unimodal tracking, experiments are conducted independently on the RGB and IR modalities. The results are summarised in Table II. For fairness, all Tracking by Detection methods share the same detector weights, while Tracking by Query methods were retrained following their default configurations.
The performance is evaluated using standard MOT metrics [mota, hota], including MOTA, IDF1, HOTA, and ID switch. With the deformable convolution-based alignment, our solution achieves 63.26 MOTA, 56.15 IDF1, 74.92 HOTA, and 159 IDs on the RGB modality, and 79.48, 66.57, 83.75, and 131 on the IR modality. With the STN-based alignment, the results were 63.08, 56.43, 74.99, and 151 on RGB, and 79.91, 66.81, 84.62, and 138 on IR. These results demonstrate that the proposed multimodal fusion method consistently outperforms existing unimodal approaches across all metrics, with particularly notable improvements in HOTA and IDF1. For example, compared with the strongest unimodal baseline MOTIP, the deformable alignment strategy improves HOTA and IDF1 by 3.11% and 9.86% on the RGB modality, while the STN-based strategy achieves gains of 3.39% and 9.93%, respectively.
@@ Acknowledgement
[Uncaptioned image] Tianyang Xu (Member, IEEE) received the B.Sc. degree in electronic science and engineering from Nanjing University, Nanjing, China, in 2011. He received his PhD degree at the School of Artificial Intelligence and Computer Science, Jiangnan University, Wuxi, China, in 2019. He was a Research Fellow in the Centre for Vision, Speech and Signal Processing (CVSSP), University of Surrey. He is currently an Associate Professor at the School of Artificial Intelligence and Computer Science, Jiangnan University, Wuxi, China. His research interests include visual understanding and multi-modal learning. He has published several scientific papers, including IEEE TPAMI, IJCV, IEEE TIP, CVPR, ICCV, ICML, NeurIPS, etc. His publications have been cited more than 6000 times. He achieved top 1 performance in several competitions, including the VOT2020 RGBT challenge, Anti-UAV challenge, and MMVRAC challenges.
[Uncaptioned image] Jinjie Gu is currently pursuing his Master’s degree at Jiangnan University. His research focuses on object detection and multi-modal visual object tracking, including single-object tracking, multi-object tracking, and anti-UAV tracking. He served as one of the organisers of the 4th Anti-UAV Workshop at CVPR2025. He participated in several challenges on countering maneuvering targets, achieving top-two rankings in the multi-object tracking track of the PETS2025 Challenge.
[Uncaptioned image] Xuefeng Zhu received the B.Eng. degree in Internet of Things Engineering from Chongqing University of Posts and Telecommunications, Chongqing, China, in 2017. He received the PhD degree in pattern recognition and intelligent system, from Jiangnan University, Wuxi, China, in 2024. He was a visiting PhD student with the Centre for Vision, Speech and Signal Processing (CVSSP), University of Surrey, Guildford, U.K. from 2022 to 2023. Currently, he is a lecturer with the School of Artificial Intelligence and Computer Science, Jiangnan University, Wuxi, China. His research interests include visual object tracking and multi-modal learning. He has published several scientific papers, including IJCV, IEEE TMM, IEEE TCSVT, PR, AAAI, etc. He achieved the 1st place award in the VOT2020 RGB-T challenge (ECCV2020).
[Uncaptioned image] Xiao-Jun Wu received the BS degree in mathematics from Nanjing Normal University, Nanjing, PR China, in 1991, the MS degree from the Nanjing Univer-sity of Science and Technology, Nanjing, China, in 1996, and the PhD degree in pattern recognition and intelligent system, from the Nanjing University of Science and Technology, in 2002. He was a fellow of United Nations University, International Institute for Software Technology (UNU/IIST) from 1999 to 2000. He joined Jiangnan University in 2006 where he is currently a distinguished professor with the School of Artificial Intelligence and Computer Science, Jiangnan University. He has published more than 400 papers in his fields of research. He was a visiting postdoctoral Researcher in the Centre for Vision, Speech, and Signal Processing (CVSSP), University of Surrey, U.K. from 2003 to 2004, under the supervision of Professor Josef Kittler. His current research interests are pattern recognition, computer vision, fuzzy systems, and neural networks. He owned several domestic and international awards because of his research achievements. Currently, he is a fellow of IAPR and AAIA.
[Uncaptioned image] Josef Kittler (M’74-LM’12) received the B.A., Ph.D., and D.Sc. degrees from the University of Cambridge, in 1971, 1974, and 1991, respectively. He is a distinguished Professor of Machine Intelligence at the Centre for Vision, Speech and Signal Processing, University of Surrey, Guildford, U.K. He conducts research in biometrics, video and image database retrieval, medical image analysis, and cognitive vision. He published the textbook Pattern Recognition: A Statistical Approach and over 700 scientific papers. His publications have been cited more than 78,000 times (Google Scholar). He is series editor of Springer Lecture Notes on Computer Science. He currently serves on the Editorial Boards of Pattern Recognition Letters, Pattern Recognition and Artificial Intelligence, Pattern Analysis and Applications. He also served as a member of the Editorial Board of IEEE Transactions on Pattern Analysis and Machine Intelligence during 1982-1985. He served on the Governing Board of the International Association for Pattern Recognition (IAPR) as one of the two British representatives during the period 1982-2005, President of the IAPR during 1994-1996.

### v2

@@ II-B UAV-Targeted Tracking Benchmarks
For single-object tracking, Chen et al. [12] introduced USC Drone, a dataset with 60 sequences and 20K frames, accompanied by a Fast-RCNN + MDNet system where the residual image input improves adaptation to a fast UAV motion. Huang et al. [25] constructed Anti-UAV410, a large-scale thermal infrared benchmark with 410 sequences and 438K frames, and proposed SiamDT with dual-semantic features for tiny targets. Zhu et al. [70] released Anti-UAV600, comprising 600 thermal sequences and 723K frames without prior bounding-box initialisation, and designed a “global detection + local tracking” framework that avoids reliance on the initial frame information through evidential collaboration. Despite the progress, these benchmarks remain limited to SOT and cannot depict scenarios involving multiple UAVs.
For multi-UAV tracking, UAVSwarm [49] builds a visible-light dataset with 72 sequences (12K frames) and validates a Faster-RCNN/YOLOX + ByteTrack baseline. MOT-FLY [14] presents an air-to-air high-resolution dataset (1920×1080) with 16 sequences (11K frames) and benchmarked advanced MOT frameworks. Recently, the 4th Anti-UAV Challenge [1] released the first large-scale infrared multi-UAV benchmark with 300 sequences and 225K frames.
@@ II-C Multi-Modal Tracking Potential
The performance bottleneck exhibited by a single visual modality in complex scenarios have driven the development of multi-modal fusion technologies [13], such as RGB-T (RGB + thermal infrared), RGB-D (RGB + depth), and RGB-E (RGB + event). Integrating the complementary information conveyed by different modalities significantly enhances the tracking robustness.
@@
However, multi-modal research in anti-UAV tasks remains scarce. The exceptions include the Anti-UAV dataset [28], which was the first to introduce an RGB-T dataset, with 348 sequences and 297K frames. As a baseline, [28] proposed a dual-stream feature semantic consistency training strategy. Yet, due to the modalities being unaligned, it cannot exploit the real potential of the cross-modal complementarity, instead using them independently. Similarly, the Halmstad Drone dataset [46] includes 650 videos (365 for IR and 285 for RGB) along with 90 audio clips, but focuses on building automated multi-sensor systems rather than investigating deep learning-based fusion. Its incomplete correspondence between IR and visible-light data also limits the utility of multi-modal anti-UAV research. Moreover, these datasets primarily target detection or SOT, leaving a void in multi-modal multi-UAV tracking benchmarks and methods.
@@
A fundamental challenge in multi-modal tracking lies in modality misalignment. While most studies assume the data being spatio-temporally aligned, real-world data often suffers from misalignment due to differences in imaging principles, field-of-view, resolution, and transmission delays. This problem is especially critical for UAV tracking, where the targets are small and fast-moving, and even slight offsets undermine effective fusion. Traditional pixel-level solutions rely on feature-point matching and affine/perspective transformations to achieve alignment, but these methods incur considerable preprocessing latency and are unsuitable for real-time multi-UAV tracking. Recent studies, therefore, emphasise learnable feature alignment. [61] proposes an aligned region CNN to align RGB and thermal regions using a Region Feature Alignment Module with RoI jitter. [9] proposes cross-modality spatial offset modelling, together with an offset-guided deformable alignment and fusion, achieving robust UAV RGB-IR detection without strict pre-alignment. Inspired by these advances, we adopt two feature-level alignment strategies for UAV tracking: (1) learning sampling offsets via deformable convolution [17]; and (2) predicting affine transformation matrices via a spatial transformer network [27]. Both strategies are end-to-end trainable without additional losses, enabling efficient mitigation of misalignment in multi-modal fusion.
@@
Additionally, given the small size and irregular motion of UAVs, as well as the inability of traditional RGB or IR modalities to distinguish targets from background under similar appearance or temperature, this study further introduces the Event modality. As a cutting-edge visual sensor, event cameras capture asynchronous event streams by recording changes in light intensity. They offer unique advantages such as high dynamic range and microsecond-level temporal resolution, making them particularly suited for extreme scenarios. Existing event-based tracking research has produced datasets such as VisEvent [51], FE108 [60], EventVOT [52], and COESOT [47], along with methods like CEUTrack [47], TENet [42], ViPT [69], and AFNet [59], but are restricted to the single-object tracking task, leaving a critical gap in event-based multi-UAV tracking.
@@ III MM-UAV BENCHMARK
The absence of a dedicated multi-modal multi-UAV tracking dataset to support anti-UAV research has significantly impeded the study of multi-modal fusion technologies. To address this, this paper presents the first large-scale tri-modal multi-UAV tracking dataset, with each modality containing 1321 sequences (2.8 million frames), aiming to fill this critical research void. This section gives a comprehensive overview of the dataset, covering core aspects such as data acquisition, annotation protocols, statistical distributions, and challenging attributes.
@@ IV-D MMA-SORT
Existing multi-object tracking models (such as SORT [53] and its variants [40, 37, 6, 62, 2, 43, 44]) exhibit notable limitations in UAV-target tracking tasks. First, the small size and complex motion patterns of UAV targets pose a significant challenge to motion modelling. The linear motion assumption inherent to the Kalman filter often leads to substantial discrepancies between the predicted bounding boxes (based on previous frames) and the actual detections in the current frame. This renders trackers centred on IOU metrics ineffective in achieving cross-frame identity association, thereby resulting in a large number of ID switches (IDSW). To address this issue, we downplay the role of IOU metrics in the process of target identity association. Unlike traditional trackers that treat IOU matching as an essential prerequisite for identity association, we argue that Kalman filters are less reliable in modelling the motion of small, irregularly moving UAV targets, compared to targets like pedestrians and vehicles. Thus, rigidly adhering to IOU matching as a prerequisite in multi-UAV tracking tasks is overly restrictive.
@@
where $\mathbf{d}$ represents the final distance, $d_{\text{iou}}$ is the IOU distance between the detection box and the trajectory’s predicted box, and $d_{\text{app}}$ is the appearance embedding distance between the detection and the trajectory. The calculation for IOU distance and appearance distance follows those in BoT-SORT [2].
@@ V-A Implementation Details
In the proposed framework, the detector’s backbone, the FPN module, and the detection head follow the original YOLOX [23]. For RGB images (640×360) and IR images (640×512) from the two modalities, we first resize them to 640×640 using the standard YOLOX preprocessing pipeline before feeding them into the backbone for feature extraction. The proposed OGAA and ADFM modules are inserted between the backbone and FPN. They are responsible for cross-modal alignment and fusion of the three-scale feature maps extracted independently by the two modality-specific backbones. The fused multi-scale features are then forwarded to the respective FPN modules for subsequent detection, and the final detection results are passed to the MMA-SORT tracker for multi-object tracking.
@@
For the appearance-based ReID module, given the limited visual information of UAV targets, we directly adopt the ReID model and training strategy of DeepSORT [40], using a ResNet backbone to extract discriminative appearance features.
During inference, the detection confidence threshold is 0.3. For the tracker, the IoU threshold $\theta_{\text{iou}}$ is 0.9, the appearance threshold is 0.2, the motion threshold $\theta_{\text{motion}}$ is 0.2, and the confidence threshold for initialising new tracks is 0.7.
@@ V-B Quantitative Comparison
For comparative evaluation, two major categories of state-of-the-art multi-object tracking methods were systematically evaluated, including Tracking by Detection approaches (DeepSORT [40], BoT-SORT [2], OC-SORT [6], Deep OC-SORT [37], BoostTrack [43], BoostTrack++ [44]) and Tracking by Query approaches (TransTrack [45], MOTR [58], MOTRv2 [63], MeMOTR [22], MOTIP [21]). All trackers are retrained and tested on our MM-UAV to enable a fair comparison. Since existing methods are designed for unimodal tracking, experiments are conducted independently on the RGB and IR modalities. The results are summarised in Table II. For fairness, all Tracking by Detection methods share the same detector weights, while Tracking by Query methods were retrained following their default configurations.
The performance is evaluated using standard MOT metrics [5, 36], including MOTA, IDF1, HOTA, and ID switch. With the deformable convolution-based alignment, our solution achieves 63.26 MOTA, 56.15 IDF1, 74.92 HOTA, and 159 IDs on the RGB modality, and 79.48, 66.57, 83.75, and 131 on the IR modality. With the STN-based alignment, the results were 63.08, 56.43, 74.99, and 151 on RGB, and 79.91, 66.81, 84.62, and 138 on IR. These results demonstrate that the proposed multimodal fusion method consistently outperforms existing unimodal approaches across all metrics, with particularly notable improvements in HOTA and IDF1. For example, compared with the strongest unimodal baseline MOTIP, the deformable alignment strategy improves HOTA and IDF1 by 3.11% and 9.86% on the RGB modality, while the STN-based strategy achieves gains of 3.39% and 9.93%, respectively.